OpenStudy (kva1992):

How would one find the solution for 4cosx-2secx=0? I have a feeling I'm over thinking the problem.

OpenStudy (kva1992):

furthest im getting is 4cosx -1/2cosx = 0

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ 4\cos x-2\sec x=0 }\) \(\LARGE\color{blue}{ 2(2\cos x-\sec x)=0 }\) \(\LARGE\color{blue}{ 2\cos x-\sec x=0 }\) ──────────────────────── \(\LARGE\color{blue}{ 2\cos x-\frac{1}{\cos x}=0 }\) everything times `cos x` \(\LARGE\color{blue}{ 2\cos^2x-1=0 }\) \(\LARGE\color{blue}{ 2\cos^2x=1 }\) \(\LARGE\color{blue}{ \cos^2x=1/2 }\) \(\LARGE\color{blue}{ \cos x=±\sqrt{1/2} }\)

OpenStudy (solomonzelman):

That would be π /4

OpenStudy (kva1992):

alright i have a question now

OpenStudy (kva1992):

what happened to the factored 2 does it just disappear?